OpenStudy (anonymous):

1. A health department estimates that 8 vials of malaria serum will treat 100 people. At this rate, how many vials are required to treat 175 people? Write your answer as a complete sentence

OpenStudy (texaschic101):

This is a proportion.. 8/100 = x/175 --- 8 vials to 100 people = x vials to 175 people now we cross multiply (100)(x) = (8)(175) 100x = 1400 -- now divide both sides by 100 x = 14 If 8 vials of malaria serum treats 100 people, then 14 vials treat 175 people.